# A complex-projected Rayleigh quotient iteration for targeting interior eigenvalues

### v2

@@ 1 Introduction
In this paper, we propose a new algorithm that modifies RQI to overcome this problem by exploiting any prior information known about the target eigenvector. The idea of the new method is as follows. Starting with an approximation of the target eigenvector, we perturb the original problem by adding a purely complex projection matrix constructed using this approximation, such that the unwanted eigenvalues are lifted into the complex plane, while the target eigenvalue remains close to the real line (see Figure 1 in Section 3), thereby artificially increasing the gap between the target eigenvalue and its neighbours. We then use the perturbed matrix in RQI, where in each step of the iteration we update the perturbation matrix using the previous eigenvector approximation and gradually reduce the size of the perturbation.
@@
$A\longrightarrow A+i\gamma(I-\bm{u}\bm{u}^{\ast})\,,$
where $i$ denotes the imaginary unit, the asterisk denotes the complex conjugate transpose and $I-\bm{u}\bm{u}^{*}$ is the projection onto the orthogonal complement of $\bm{u}$. The real scalar $\gamma>0$ is used to control the magnitude of the perturbation.
@@
approximately. Since $\mu\in\mathbb{R}$, because $A$ is Hermitian, the perturbed eigenvalues now lie (approximately) along the line $i\gamma$ in the upper half of the complex plane, again see Figure 1, and so the distance between the perturbed target eigenvalue $\widetilde{\lambda}$ and the remainder of the perturbed spectrum is approximately $\gamma$. Applying RQI to this problem, we will not compute the target eigenvalue but rather an eigenvalue of the perturbed matrix. Therefore, we replace $\gamma$ by a non-negative decreasing zero sequence ${(\gamma^{(k)})}_{k\geq 0}\subset\mathbb{R}$ such that the perturbation becomes increasingly small as $k\rightarrow\infty$ and we recover the original unperturbed problem. If the sequence ${(\gamma^{(k)})}_{k\geq 0}$ is chosen to decay sufficiently fast, then asymptotically our algorithm converges like classic RQI either cubically or quadratically to an eigenpair.
@@
The remainder of the paper is organised as follows. In the next section, we review convergence results for Inverse Iteration and classic RQI, since they form the basis for the convergence analysis of our method. In Section 3, we then introduce the new Projected Rayleigh Quotient Iteration (PRQI) method. We also prove there that the approach introduced above is equivalent to perturbing the matrix according to
@@
That is, it is not necessary to perturb $A$ by the full projection matrix $i\gamma(I-\bm{u}\bm{u}^{\ast})$ but it suffices to shift only the diagonal entries by $i\gamma$. This simplifies the implementation and also allows to prove asymptotic convergence for our method. Finally, in Section 4 we present three numerical examples that demonstrate the efficacy of our method, including a practical example from an application in photonics. In particular, we present results where our algorithm succeeds in locating the target eigenvalue within a closely-spaced spectrum. Often a crude approximation is sufficient for our method to succeed in converging to the target eigenpair, while for the same initial conditions classic RQI fails.
@@ 2 Classic Rayleigh Quotient Iteration
[…]
@@ 3 Projected Rayleigh Quotient Iteration
In this section, we present the new Projected Raleigh Quotient Iteration (PRQI), along with a simplified version of the algorithm and an analysis of the local convergence properties. First, we provide some intuition behind the main idea of introducing a complex projection into the shift.
@@
i.e., the eigenvalues of $A$ (except for the target eigenvalue $\lambda_{1}$) are raised into the complex plane and the distance between the target eigenvalue and the remaining eigenvalues is now approximately $\gamma$, see Figure 1.
Figure 1: Section of the spectrum of a matrix $A$ (orange circles) and the perturbed matrix $\widetilde{A}$ (blue triangles) for $\gamma=1$. Most of the perturbed eigenvalues $\widetilde{\lambda}$ have $\text{Im}({\widetilde{\lambda}})\approx 1$ while one eigenvalue lies close to the real line. The matrix $A$ and the vector $\bm{u}$ that were used to construct $\widetilde{A}$ for this plot come from Example 3 in Section 4.
As remarked in the introduction, the basins of attractions of RQI depend critically on the eigenvalue gap. By increasing the eigenvalue gap we expect the basin of attraction around the target eigenvector to increase, thus making the algorithm more likely to succeed.
Of course, by applying RQI to the matrix $\widetilde{A}$ defined in (3.1), we would compute an eigenpair of $\widetilde{A}$ instead of $A$. Therefore, we replace the fixed shift $\gamma$ by a non-negative sequence ${(\gamma^{(k)})}_{k\geq 0}$ that converges to 0 and we replace the initial approximation $\bm{u}$ by the current vector iterate $\bm{x}^{(k)}$. The full iteration is summarised in Algorithm 3.1. The stopping criteria can be chosen to be the same as for classic RQI, e.g., terminating when either the norm of the residual (2.5) is below a given tolerance or a maximum number of iterations is reached.
@@
Note that in practice we allow each $\gamma^{(k)}$ to depend on the previous iterates $\bm{x}^{(0)},$ $\bm{x}^{(1)},\ldots,\,\bm{x}^{(k)}$. Hence, the sequence ${(\gamma^{(k)})}_{k\geq 0}$ is not an input to Algorithm 3.1, but instead we have included Step 3 where $\gamma^{(k)}$ is updated. How the sequence ${(\gamma^{(k)})}_{k\geq 0}$ should be chosen to ensure both stability of the iteration and rapid convergence, is discussed in Section 3.2. First, we derive a simplified version of the algorithm.
@@ 3.1 A simplified algorithm
[…]
@@
Now, observe that $C=B-i\gamma\bm{u}\bm{u}^{\ast}$. The Sherman–Morrison formula (see, e.g., [12, p. 50]) states that $C$ is invertible if $i\gamma\bm{u}^{\ast}B^{-1}\bm{u}\neq-1$. To check that this holds, let $A=VDV^{\ast}$ be an eigendecomposition of $A$ with a unitary matrix $V$, whose columns are the eigenvectors of $A$, and a diagonal matrix $D=\mathrm{diag}(\lambda_{1},\dotsc,\lambda_{n})$ containing the corresponding eigenvalues. We then have $B=V(D-\mu I+i\gamma I)V^{\ast}$ and consequently
@@
$\bm{u}^{\ast}B^{-1}\bm{u}$ $=\bm{u}^{\ast}V{(D-\mu I+i\gamma I)}^{-1}V^{\ast}\bm{u}$
$=\bm{w}^{\ast}{(D-\mu I+i\gamma I)}^{-1}\bm{w}\qquad=\sum_{j=1}^{n}\frac{\left\lvert w_{j}\right\rvert^{2}}{\lambda_{j}-\mu+i\gamma}\,,$
where we introduced $\bm{w}\coloneqq V^{\ast}\bm{u}$. Multiplying by $i\gamma$ and taking the real part gives
[…]
In particular, $i\gamma\bm{u}^{\ast}B^{-1}\bm{u}\neq-1$ so that $C$ is also invertible.
Therefore, with $\alpha\coloneqq 1+i\gamma\bm{u}^{\ast}B^{-1}\bm{u}\neq 0$, the Sherman–Morrison formula gives
$C^{-1}\bm{u}$ $={\big{(}B+i\gamma\bm{u}\bm{u}^{\ast}\big{)}}^{-1}\bm{u}$
$=\big{(}B^{-1}-\alpha^{-1}B^{-1}\bm{u}i\gamma\bm{u}^{\ast}B^{-1}\big{)}\bm{u}$
$=B^{-1}\bm{u}-B^{-1}\bm{u}\underbrace{\alpha^{-1}i\gamma\bm{u}^{\ast}B^{-1}\bm{u}}_{\in\mathbb{C}}=B^{-1}\bm{u}(1-\alpha^{-1}i\gamma\bm{u}^{\ast}B^{-1}\bm{u})\,.$
Thus, the vector $C^{-1}\bm{u}$ is a complex scalar multiple of $B^{-1}\bm{u}$. ∎
@@ 3.2 Convergence
We now turn to the question of how to choose ${(\gamma^{(k)})}_{k\geq 0}$. Throughout this section, $(\lambda_{1},\bm{v}_{1})$ denotes the target eigenpair of $A$ and ${(\bm{x}^{(k)})}_{k\geq 0}$ is the sequence of vectors that is obtained by Algorithm 3.2. When choosing ${(\gamma^{(k)})}_{k\geq 0}$ there are two things that should be balanced: the effect of the perturbation on the gap between the target eigenvalue and the remainder of the spectrum, and the rate of convergence of the iterates to the target. If $\gamma^{(k)}$ decays too rapidly, then we lose the effect of the perturbation and the algorithm will essentially behave like the classic RQI. On the other hand, if $\gamma^{(k)}$ decays too slowly then the iterates $\bm{x}^{(k)}$ will stay close to an eigenvector of the perturbed matrix, thus degrading the overall convergence. To achieve this balance, we will choose $\gamma^{(k)}$ depending on the current iterate $\bm{x}^{(k)}$ and decaying at the same rate as the expected convergence rate.
@@
which leads to cubic convergence as in Theorem 2.5. In light of this, we will assume that $\gamma^{(k)}$ is chosen to decay either linearly or quadratically in $\tan(\theta(\bm{x}^{(k)},\bm{v}_{1})$.
@@
As we will show in the following theorem, this results in an asymptotic convergence rate that is either quadratic (if $q=0$ in (3.3)) or cubic (if $q=1$ in (3.3)). In practice, we have observed that choosing $\gamma^{(k)}$ to obtain quadratic convergence ($q=0$ in (3.3)) usually only leads to a small number of extra iterations to achieve the desired error tolerance, but has the benefit of adding stability to the method.
@@
Remark.
The result is a special case of [8, Thm. 2.2], which discusses the more general case of shifted Inverse Iteration with variable shifts $\sigma^{(k)}$ and inexact solves of the linear system $(A-\sigma^{(k)}I)\bm{y}^{(k+1)}=\bm{x}^{(k)}$, i.e., the linear system is only solved up to a tolerance $\tau^{(k)}\geq 0$ so that $\|{(A-\sigma^{(k)}I)\bm{y}^{(k+1)}-\bm{x}^{(k)}}\|\leq\tau^{(k)}$. Our case corresponds to the choice of $\sigma^{(k)}=\mu^{(k)}-i\gamma^{(k)}$ and $\tau^{(k)}=0$, for all $k$. Since the proof in [8] is only given for real symmetric $A$, we include a direct proof of our case for completeness.
@@
We interpret one step of Algorithm 3.2 as one step of Inverse Iteration applied to the matrix $A$ using the shift $\sigma^{(0)}\coloneqq\mu^{(0)}-i\gamma^{(0)}$. Since for all complex vectors $\bm{z}\in\mathbb{C}^{n}$ and all Hermitian matrices $A=A^{\ast}$ we have $\mathcal{R}_{A}(\bm{z})\in\mathbb{R}$ (which can be easily seen by writing $\bm{z}=\bm{a}+i\bm{b}$ and expanding the Rayleigh Quotient of $\bm{z}$) and thus $\mu^{(0)}\in\mathbb{R}$, Equation (3.4) implies that
$|\lambda_{1}-\sigma^{(0)}|<|\lambda_{2}-\sigma^{(0)}|\leq\cdots\leq|\lambda_{n}-\sigma^{(0)}|\,.$
We apply the convergence theorem of Inverse Iteration (Theorem 2.1) to obtain
@@
Let us first bound the numerator in the right hand side. By using the quadratic accuracy of the Rayleigh Quotient (2.4) to bound $|\lambda_{1}-\mu^{(0)}|$, recalling that $\mu^{(0)}=\mathcal{R}_{A}(\bm{x}^{(0)})$, and the assumption (3.3) to bound $\gamma^{(0)}$ we have
$|\lambda_{1}-\sigma^{(0)}|$ $\leq|\lambda_{1}-\mu^{(0)}|+|\gamma^{(0)}|$
$\leq\|A-\lambda_{1}I\|\tan^{2}(\theta(\bm{x}^{(0)},\bm{v}_{1}))+C_{\gamma}\tan^{1+q}(\theta(\bm{x}^{(0)},\bm{v}_{1}))$
$\leq\big{(}\|A-\lambda_{1}I\|+C_{\gamma}\big{)}\tan^{1+q}(\theta(\bm{x}^{(0)},\bm{v}_{1}))\,.$ (3.8)
In the last step, we have used that (3.5) implies $\tan(\theta(\bm{x}^{(0)},\bm{v}_{1}))\leq c<1$, since
$\|A-\lambda_{1}I\|=\max_{k=2,3,\ldots,n}|\lambda_{k}-\lambda_{1}|\geq|\lambda_{2}-\lambda_{1}|$ (3.9)
and $C_{\gamma}>0$. Hence, $\tan^{2}(\theta(\bm{x}^{(0)},\bm{v}_{1}))\leq\tan^{1+q}(\theta(\bm{x}^{(0)},\bm{v}_{1}))$ for both $q=0$ and $q=1$.
@@
To compare (2.7) and (3.5), using again (3.9) and $C_{\gamma}>0$ gives
@@
This is not very surprising since we essentially apply RQI to the matrix $A$ but then use the “wrong” matrix $A-i\gamma^{(k)}I$ when solving the linear system during the iteration. Also, as described at the beginning of Section 3, the perturbation at each step (cf. (3.1)) increases the gap between the target eigenvalue and remainder of the spectrum, which may allow us to weaken the condition (3.4). However, this property has not been properly quantified and as such, has not been used in the proof of our convergence result. We therefore believe that in order to show that our method does allow for less accurate initial guesses, the original formulation as given in Algorithm 3.1 has to be considered and the effect of the perturbation (cf. (3.1)) has on the spectrum needs to be quantified.
@@
This shows that $\gamma^{(k)}=\left\lVert\bm{r}^{(k)}\right\rVert^{1+q}$ satisfies Assumption 1 for $q\in\{0,1\}$ and hence, Theorem 3.2 holds with $C_{\gamma}=\operatorname{spread}(A)^{1+q}$.

### v3

@@ Abstract
Key words. eigenvalues, Rayleigh quotient iteration, complex shifts, eigenvector information, localised eigenvectors, larger convergence radius
MSC codes. 65F15, 65L15, 65N25
@@ 1 Introduction
In this paper, we propose a new algorithm that modifies RQI to overcome this problem by exploiting any prior information known about the target eigenvector. The idea of the new method is as follows. Starting with an approximation of the target eigenvector, we perturb the original problem by adding a purely complex projection matrix constructed using this approximation, such that the unwanted eigenvalues are lifted into the complex plane, while the target eigenvalue remains close to the real line (see Figure 1), thereby artificially increasing the gap between the target eigenvalue and its neighbours. We then use the perturbed matrix in RQI, where in each step of the iteration we update the perturbation matrix using the previous eigenvector approximation and gradually reduce the size of the perturbation.
Figure 1: Section of the spectrum of a matrix $A$ (orange circles) and the perturbed matrix $\widetilde{A}$ (blue triangles). Most of the perturbed eigenvalues $\widetilde{\lambda}$ have $\text{Im}({\widetilde{\lambda}})\approx 1$ while one eigenvalue lies close to the real line. The matrix $A$ and the vector $\bm{u}$ that were used to construct $\widetilde{A}$ for this plot come from Example 3 in Section 4.
@@
$A\longrightarrow A+i\gamma(I-\bm{u}\bm{u}^{\ast})\,,$ (1.1)
where $i$ denotes the imaginary unit, the asterisk denotes the complex conjugate transpose and $I-\bm{u}\bm{u}^{\ast}$ is the projection onto the orthogonal complement of $\bm{u}$. The real scalar $\gamma>0$ is used to control the magnitude of the perturbation.
@@
approximately. Since $\mu\in\mathbb{R}$, because $A$ is Hermitian, the perturbed eigenvalues now lie (approximately) along the line $i\gamma$ in the upper half of the complex plane, again see Figure 1, and so the distance between the perturbed target eigenvalue $\widetilde{\lambda}$ and the remainder of the perturbed spectrum is approximately $\gamma$. Applying RQI to this problem, we will not compute the target eigenvalue but rather an eigenvalue of the perturbed matrix. Therefore, we replace $\gamma$ by a non-negative decreasing zero sequence ${(\gamma^{(k)})}_{k\geq 0}\subset\mathbb{R}$ such that the perturbation becomes increasingly small as $k\rightarrow\infty$ and we recover the original unperturbed problem.
In practice, the size of the perturbation $\gamma^{(k)}$ needs to be balanced. If it decays too fast then we lose the effect of the complex perturbation and our method may not converge. On the other hand, if it decays too slowly then it will slow down the convergence to the target eigenpair. A good balance is to choose $\gamma^{(k)}$ at each step to be the norm (or norm squared) of the current residual. In this way, our algorithm utilises approximate eigenvector information to choose both the direction as well as the magnitude of the complex perturbation. We prove that this choice leads to an asymptotic rate of convergence that is either quadratic or cubic, depending on the choice of the perturbations $\gamma^{(k)}$.
@@
The remainder of the paper is organised as follows. In the next section, we review convergence results for Inverse Iteration and classic RQI, since they form the basis for the convergence analysis of our method. In Section 3, we then introduce the new Projected Rayleigh Quotient Iteration (PRQI) method. We also prove there that the approach introduced above in (1.1) is equivalent to perturbing the matrix according to
@@
That is, it is not necessary to perturb $A$ by the full projection matrix $i\gamma(I-\bm{u}\bm{u}^{\ast})$ as in (1.1) but it suffices to shift only the diagonal entries by $i\gamma$. This simplifies the implementation and also allows us to prove asymptotic convergence of our method in Section 3.2. Finally, in Section 4 we present three numerical examples that demonstrate the efficacy of our method, including a practical example from an application in photonics. In particular, we present results where our algorithm succeeds in locating the target eigenvalue within a closely-spaced spectrum. Often a crude approximation is sufficient for our method to succeed in converging to the target eigenpair, while for the same initial conditions classic RQI fails.
@@ 2 Classic Rayleigh Quotient Iteration
$|\mu-\lambda|\leq\frac{\left\lVert\bm{r}\right\rVert^{2}}{\delta}\qquad\text{and}\qquad\sin\theta(\bm{u},\bm{v})\leq\frac{\left\lVert\bm{r}\right\rVert}{\delta}\,.$
@@ 3 Projected Rayleigh Quotient Iteration
In this section, we present the new Projected Rayleigh Quotient Iteration (PRQI), along with a simplified version of the algorithm and an analysis of the convergence properties. First, we provide some intuition behind the main idea of introducing a complex projection into the shift.
@@
i.e., the eigenvalues of $A$ (except for the target eigenvalue $\lambda_{1}$) are raised into the complex plane and the distance between the target eigenvalue and the remaining eigenvalues is now approximately $\gamma$ (again, see Figure 1).
By applying RQI to the matrix $\widetilde{A}$ defined in (3.1), we would compute an eigenpair of $\widetilde{A}$ instead of $A$. Therefore, we replace the initial approximation $\bm{u}$ by the current vector iterate $\bm{x}^{(k)}$ and use the resulting matrix in a shift-invert method with the shift chosen as the Rayleigh quotient of $\bm{x}^{(k)}$ with respect to $A$. Since the approximations in (3.2) become exact if $\bm{u}=\bm{v}_{1}$, for sufficiently accurate initial shifts the method will converge to the correct eigenpair $(\lambda_{1},\bm{v}_{1})$. Due to the increased eigenvalue gap the basin of attraction around the target eigenvector increases, thus making the algorithm more likely to succeed. To further accelerate the method, we also replace the fixed scaling parameter $\gamma$ by a non-negative sequence ${(\gamma^{(k)})}_{k\geq 0}$ that converges to 0.
The full iteration is summarised in Algorithm 3.1. The stopping criteria can be chosen to be the same as for classic RQI, e.g., terminating when either the norm of the residual (2.5) is below a given tolerance or a maximum number of iterations is reached.
@@
Note that in practice we allow each $\gamma^{(k)}$ to depend on the previous iterates $\bm{x}^{(0)},$ $\bm{x}^{(1)},\ldots,\,\bm{x}^{(k)}$. Hence, the sequence ${(\gamma^{(k)})}_{k\geq 0}$ is not an input to Algorithm 3.1, but instead we have included Step 3 where $\gamma^{(k)}$ is updated. How the sequence ${(\gamma^{(k)})}_{k\geq 0}$ should be chosen to ensure both stability of the iteration and rapid convergence is discussed in Section 3.2. First, we derive a simplified version of the algorithm.
@@ 3.1 A simplified algorithm
$B\coloneqq A-\mu I+i\gamma I\quad\text{and}\quad C\coloneqq{A-\mu I+i\gamma(I-\bm{u}\bm{u}^{\ast})}\,,$
@@
Now, observe that $C=B-i\gamma\bm{u}\bm{u}^{\ast}$. The Sherman–Morrison formula (see, e.g., [12, p. 50]) states that $C$ is invertible if $i\gamma\bm{u}^{\ast}B^{-1}\bm{u}\neq 1$. To check that this holds, let $A=VDV^{\ast}$ be an eigendecomposition of $A$ with a unitary matrix $V$, whose columns are the eigenvectors of $A$, and a diagonal matrix $D=\mathrm{diag}(\lambda_{1},\dotsc,\lambda_{n})$ containing the corresponding eigenvalues. We then have $B=V(D-\mu I+i\gamma I)V^{\ast}$ and consequently
@@
$\bm{u}^{\ast}B^{-1}\bm{u}=\bm{u}^{\ast}V{(D-\mu I+i\gamma I)}^{-1}V^{\ast}\bm{u}=\sum_{j=1}^{n}\frac{\left\lvert w_{j}\right\rvert^{2}}{\lambda_{j}-\mu+i\gamma}\,,$
where we introduced $\bm{w}\coloneqq V^{\ast}\bm{u}$. Multiplying by $i\gamma$ and taking the imaginary part gives
$\operatorname{Im}(i\gamma\bm{u}^{\ast}B^{-1}\bm{u})=\sum_{j=1}^{n}\frac{\gamma(\lambda_{j}-\mu)\left\lvert w_{j}\right\rvert^{2}}{{(\lambda_{j}-\mu)}^{2}+\gamma^{2}}\neq 0.$
In particular, $i\gamma\bm{u}^{\ast}B^{-1}\bm{u}\neq 1$ so that $C$ is also invertible. Defining $\beta\coloneqq 1-i\gamma\bm{u}^{\ast}B^{-1}\bm{u}\neq 0$, the Sherman–Morrison formula thus gives
$C^{-1}\bm{u}$ $={\big{(}B-i\gamma\bm{u}\bm{u}^{\ast}\big{)}}^{-1}\bm{u}$
$=\big{(}B^{-1}+\beta^{-1}B^{-1}\bm{u}i\gamma\bm{u}^{\ast}B^{-1}\big{)}\bm{u}=B^{-1}\bm{u}(1+\beta^{-1}i\gamma\bm{u}^{\ast}B^{-1}\bm{u})\,.$
@@ 3.2 Convergence
We now turn to the question of how to choose ${(\gamma^{(k)})}_{k\geq 0}$. Throughout this section, $(\lambda_{1},\bm{v}_{1})$ denotes the target eigenpair of $A$ and ${(\bm{x}^{(k)})}_{k\geq 0}$ is the sequence of vectors that is obtained by Algorithm 3.2.
@@
which leads to cubic convergence as in Theorem 2.5. In light of this, we will assume that $\gamma^{(k)}$ is chosen to decay either linearly or quadratically in $\tan(\theta(\bm{x}^{(k)},\bm{v}_{1})$. The following condition initially may seem restrictive, but it can be satisfied by the natural choice of defining $\gamma^{(k)}$ in terms of the residual (see Lemma 3.3 below).
@@
As we will show in the following theorem, this results in an asymptotic convergence rate that is either quadratic (if $q=0$ in (3.3)) or cubic (if $q=1$ in (3.3)). In practice, we have observed that choosing $\gamma^{(k)}$ to obtain quadratic convergence ($q=0$ in (3.3)) usually only leads to a small number of extra iterations to achieve the desired error tolerance, but has the benefit of adding stability to the method. Note that the theorem uses the same eigenvalue ordering as Theorem 2.1 (with $\mu=\mu^{(0)}=\mathcal{R}_{A}(\bm{x}^{(0)})$) and also as (2.6) in Theorem 2.5.
@@
We interpret one step of Algorithm 3.2 as one step of Inverse Iteration applied to the matrix $A$ using the shift $\sigma^{(0)}\coloneqq\mu^{(0)}-i\gamma^{(0)}$. Since $A$ is Hermitian we have $\mu^{(0)}=\mathcal{R}_{A}(\bm{x}^{(0)})\in\mathbb{R}$, and thus Equation (3.4) implies that
$|\lambda_{1}-\sigma^{(0)}|<|\lambda_{2}-\sigma^{(0)}|\leq\cdots\leq|\lambda_{n}-\sigma^{(0)}|\,,$
so that the convergence theorem of Inverse Iteration (Theorem 2.1) can be applied to obtain
@@
Let us first bound the numerator in the right hand side. By using the quadratic accuracy of the Rayleigh Quotient (2.4) to bound $|\lambda_{1}-\mu^{(0)}|$, recalling that $\mu^{(0)}=\mathcal{R}_{A}(\bm{x}^{(0)})$, and Assumption 1 to bound $\gamma^{(0)}$ we have
$|\lambda_{1}-\sigma^{(0)}|\leq(\|A-\lambda_{1}I\|+C_{\gamma})\tan^{1+q}(\theta(\bm{x}^{(0)},\bm{v}_{1}))\,.$ (3.8)
@@
To compare (2.7) and (3.5), note that
@@
This is not very surprising since we essentially apply RQI to the matrix $A$ but then use the “wrong” matrix $A-i\gamma^{(k)}I$ when solving the linear system during the iteration. We believe that in order to show that our method does allow for less accurate initial guesses, the original formulation as given in Algorithm 3.1 has to be considered and the effect of the perturbation (cf. (3.1)) on the spectrum needs to be quantified, which could be an interesting direction for future research. Another plausible approach is to interpret one step of PRQI as one step of an inexact inverse iteration, as in, e.g., [8], applied to the matrix $\widehat{A}^{(k)}\coloneqq A+i\gamma^{(k)}(I-\bm{v}_{1}\bm{v}_{1}^{\ast})$ with shift $\mu^{(k)}=\mathcal{R}_{A}(\bm{x}^{(k)})$ and where $\bm{v}_{1}\in\mathbb{C}^{n}$ is the target eigenvector. However, this method of proof again leads to (3.7) and so does not offer any improvement over the result above.